OpenStudy (anonymous):

Can someone help me with two problems about equations? Two of my answers were wrong! Multiply or divide as indicated. x^-8 times x^-2 I think that it could be x^10... Multiply or divide as indicated. Simplify your answer, if possible. b^3 times b^-3 I think that the answer is b^6..

hartnn (hartnn):

nopes, both incorrect.

hartnn (hartnn):

\(\huge a^x*a^y=a^{x+y}\)

OpenStudy (zzr0ck3r):

x^a*x^b = x^(a+b)

hartnn (hartnn):

so \(\huge x^{-8}*x^{-2}=x^{-8-2}=??\)

OpenStudy (zzr0ck3r):

hint: b^3*b^(-3) = b^3/b^3

hero (hero):

I'm used to interpreting negative exponents as inverses of the base times the index.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

@hartnn x^-6??

OpenStudy (zzr0ck3r):

what is -8-2?

hartnn (hartnn):

actually -8-2 = -10 so , it will be \(x^{-10}\) ok?

OpenStudy (anonymous):

@zzr0ck3r b^6?

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

what is a/a = b/b= 3/3 = d/d = smiley/smiley = b^3/b^3 = ?

OpenStudy (anonymous):

b^3

OpenStudy (zzr0ck3r):

a/a = b^3?

OpenStudy (zzr0ck3r):

what is 2/2?

OpenStudy (anonymous):

1

OpenStudy (zzr0ck3r):

what is 76584/76584?

OpenStudy (anonymous):

It's 1...

OpenStudy (zzr0ck3r):

what is a/a?

OpenStudy (anonymous):

a

OpenStudy (zzr0ck3r):

anything devided by itself is 1

OpenStudy (zzr0ck3r):

so what is b^3/b^3 ?

OpenStudy (anonymous):

1!!!

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

note: b^3*b^(-3) = b^(3+(-3)) = b^(3-3) = b^0 = 1

OpenStudy (anonymous):

x^-10